# T3: Multi-modal Tailless Triple-Flapping-Wing Robot for Efficient Aerial and Terrestrial Locomotion

### v1

@@ III-B Thrust Generation, Self-Righting and Flight Endurance
Figure 9: (a) The raw data and fitted curve for frequency and throttle, (a) The raw data and fitted curve for lift and frequency.
$f=-41.56thr^{2}+80.69thr-14.64$ (19)
@@
We also measured the endurance time of the robot in both stationary hovering mode and ground mode at a constant speed of $5\,\mathrm{cm/s}$. Experimental results show that, with a fully charged battery, the robot’s endurance time from takeoff to the point where the battery can no longer sustain hovering is approximately $6.5\,\mathrm{minutes}$. In contrast, the ground movement mode utilizes only two sets of flapping-wing modules, and the maximum vibration throttle is only half of that in hovering mode. As a result, the endurance time is significantly extended, reaching nearly four times that of the hovering mode, approximately $28\,\mathrm{minutes}$.
@@ III-C Remote Control Flight and Trajectory Tracking
Figure 10: (a) Attitude data curve of the T3 robot during flight trajectory tracking, (b) Position sampling curve from the FZMotion motion capture system.
T3 robot successfully conducted a series of flight experiments, including vertical takeoff and landing, multi-DOF flight, and obstacle-crossing trajectory tracking. As depicted in Fig. 10(a), the robot maintained exceptional tracking performance in pitch and roll angles, even as the yaw angle exhibited periodic fluctuations in the absence of active yaw control. Fig. 10(b) presents the position data feedback obtained from the FZMotion motion capture system during the trajectory tracking experiment.
@@
Figure 11: (a) Vertical takeoff and landing in altitude-hold mode, (b) Trajectory tracking experiment for obstacle-crossing task, (c) Multi-mode locamotion experiment, green, orange, and blue curves represent the crawling, flying, and landing processes, respectively.
@@ IV Conclusion
This paper presents the design, modeling, and experimentation of a three-winged, tailless flapping robot named T3, which achieves ground crawling and flight through the vibrations and flapping motion of its own wings. With just three actuators, T3 combines features like self-righting, ground crawling, multi-DOF flight, vertical takeoff and landing, and smooth mode switching. The robot employs three sets of annular symmetrically arranged flapping-wing module, with an overall weight of $37.4\,\mathrm{g}$. In flight mode, it achieves a maximum speed of $5.5\,\mathrm{m/s}$ and a flight endurance of $6.5\,\mathrm{minutes}$. In ground mode, it reaches a maximum speed of $5.4\,\mathrm{cm/s}$ and a crawling endurance of $28\,\mathrm{minutes}$. Through a meticulously designed chassis mechanism and passive elastic legs, T3 constrains and enhances the vibrations generated by the flapping modules in specific directions, marking the first instance of a miniature biomimetic robot achieving controlled ground motion solely through its own wing vibrations. These results show that the design of T3 offers a novel and feasible solution for efficient insect-inspired flying and crawling robots.

### v2

@@ III-B Thrust Generation, Self-Righting and Flight Endurance
Figure 9: (a) The raw data and fitted curve for frequency and throttle, (b) The raw data and fitted curve for lift and frequency.
$f=-41.56thr^{2}+80.69thr-14.64.$ (19)
@@
We also measured the endurance time of the robot in both stationary hovering mode and ground mode at a constant speed of $5\,\mathrm{cm/s}$. During crawling, although minor body oscillations and attitude tilts were observed due to vibration and aerodynamic lift, the robot remained stable and did not tip over. This is attributed to the relatively low and near-central placement of the center of mass, as well as the symmetric support provided by the legs. Experimental results show that, with a fully charged battery, the robot’s endurance time from takeoff to the point where the battery can no longer sustain hovering is approximately $6.5\,\mathrm{minutes}$. In contrast, the ground movement mode utilizes only two sets of flapping-wing modules, and the maximum vibration throttle is only half of that in hovering mode. As a result, the endurance time is significantly extended, reaching nearly four times that of the hovering mode, approximately $28\,\mathrm{minutes}$.
Figure 10: (a) Roll and pitch angle responses during hover stabilization. The setpoints are fixed at $0^{\circ}$ in both axes. Despite passive yaw drift, the controller maintains stable attitude tracking in roll and pitch. The root-mean-square error (RMSE) was $4.78^{\circ}$ for roll and $7.07^{\circ}$ for pitch, with maximum deviations of $16.65^{\circ}$ and $20.15^{\circ}$, respectively, (b) Position sampling curve from the FZMotion motion capture system.
Figure 11: (a) Vertical takeoff and landing in altitude-hold mode, (b) Trajectory tracking experiment for obstacle-crossing task, (c) Multi-mode locomotion experiment, green, orange, and blue curves represent the crawling, flying, and landing processes, respectively.
@@ III-C Remote Control Flight and Trajectory Tracking
The robot conducted a series of flight experiments, including vertical takeoff and landing, multi-DOF flight, and obstacle-crossing trajectory tracking.
As depicted in Fig. 10 (a), the robot was commanded to maintain a nominal hover orientation with fixed setpoints of $0^{\circ}$ in both roll and pitch. The SE(3)-based controller actively stabilized these axes while allowing passive yaw rotation due to the absence of active yaw control. The measured roll and pitch angles remained close to their reference values throughout the test, demonstrating acceptable stabilization performance in underactuated conditions. Quantitatively, the root-mean-square error (RMSE) was $4.78^{\circ}$ for roll and $7.07^{\circ}$ for pitch, with maximum absolute deviations of $16.65^{\circ}$ and $20.15^{\circ}$, respectively. The yaw drift observed in the experiment is primarily due to initial assembly imbalances. With proper assembly, the yaw drift can be minimized or even negligible. Although the yaw angle exhibited continuous drift, the position controller operates in the inertial frame and uses roll and pitch to redirect thrust, allowing the robot to follow the desired trajectory with acceptable accuracy. Fig.10 (b) presents the corresponding position feedback data collected via the FZMotion motion capture system.
@@ IV Conclusion
This paper presents the design, modeling, and experimental validation of a tailless, three-winged flapping robot capable of both flight and vibration-driven ground locomotion. Using only three actuators, the robot achieves vertical takeoff and landing, multi-DOF flight, self-righting, and seamless aerial-ground transitions. Weighing $37.4\,\mathrm{g}$, it reaches a flight speed of $5.5\,\mathrm{m/s}$ with $6.5\,\mathrm{minutes}$ of endurance, and a ground speed of $5.4\,\mathrm{cm/s}$ for up to $28\,\mathrm{minutes}$. Ground locomotion is achieved through asymmetric elastic legs and chassis geometry that direct flapping-induced vibrations into forward motion, without additional actuators. The final design enhances vibration transmission and directional bias through curved legs and tilted support structures. An SE(3)-based controller enables acceptable flight stabilization in underactuated conditions, despite the lack of yaw control. These results demonstrate a compact, efficient, and feasible approach for multi-modal locomotion in lightweight aerial-ground robots.
Future work will investigate how varying flapping frequency affects crawling speed and turning dynamics potentially enabling actuator-free control of terrestrial motion. We also plan to explore passive or minimally active yaw stabilization strategies to improve heading control and enhance robustness in underactuated flight.